OpenStudy (anonymous):

Find the coordinates of Q given that M is the midpoint of line segment PQ: P(6, -2) and M(0, 5)

OpenStudy (anonymous):

use the mid-point formula and equate corresponding x and y coordinates. u'll get, q=(-6,12)

OpenStudy (anonymous):

\[\frac{x _{1}+6}{2}=0\]\[x_{1}=(2*0)-6\]\[x_{1}=-6\]\[\frac{y _{1}+(-2)}{2}=5\]\[y _{1}=(2*5)+2\]\[y _{1}=12\]\[Q(-6,12)\] if I desrver a medal click good answer.TY

OpenStudy (anonymous):

Thanks !!!!

OpenStudy (anonymous):

alway wc